Suppose you hold a portfolio consisting of a $50,000 investment in each of 8 different common stocks. The portfolio’s beta is 1.25. Now suppose you decided to sell one of your stocks that has a beta of 1.00 and to use the proceeds to buy a replacement stock with a beta of 1.55. What would the portfolio’s new beta be?

options: 1.17, 1.23, 1.29, 1.32, 1.43
Answer #1

Increase in beta as a result of net sales=(1.55-1)=0.55

Hence increase/stock=0.55/8

=0.06875

Hence new beta=1.25+0.06875

=1.32(Approx).